OpenStudy (anonymous):

simplify square root of 28 divided by square root of 25

OpenStudy (anonymous):

\[\sqrt{28}=\sqrt{4*7}=2\sqrt{7}\] \[\sqrt{25}=?\]

OpenStudy (anonymous):

Root 28 turns into root 4 multiplied by 7 once you break them up into two terms. The denominator turns into 5 since the square root of 25 is 5. This is as far as you can go with this kind of equation.